OpenStudy (rosales0228):

Kelly tells you that when variables are in the denominator, the equation one over two plus three over x equals three over four becomes unsolvable. Kelly explains, "There is a value for x that makes the denominator zero, and you can't divide by zero." Demonstrate to Kelly how the equation is still solvable and explain your reasoning

OpenStudy (retireed):

It is not "There is a value of x that makes the denominator zero the function is unsolvable" Kelly should have said "If there is a value of x that makes the denominator 0 then the function is undefined and not solvable." 1/2 +3/x = 3/4 One equation with one unknown solve for x.

OpenStudy (dayakar):

1/2 +3/x =3/4 2/4 +3/x = 3/4 [i write 1/2 = 2/4] 3/x = 3/4 -2/4 [ transpose 2/4 ] 3/x = 1/4 x/3 = 4/1 x= 12